# Generative AI and Large Language Models for Cyber Security: All Insights You Need

### v1

@@ 80 Q 500 Q 2k Q 10k Q
V General LLMs
Tables VI, VII, VIII compare general transformer-based Large Language Models. LLM models are generally trained on a diverse and broad range of data to provide a relatively comprehensive understanding. They can handle various language tasks like translation, summarization, and question-answering. In contrast, code-specific LLMs are specialized models trained primarily on programming languages and related technical literature, which makes their primary role in understanding and generating programming code well-suited for tasks like automated code generation, code completion, and bug detection.
V-A Prevalent LLMs
V-A1 GPT-3
V-A2 GPT-4
V-A3 T5
V-A4 BERT
V-A5 ALBERT
V-A6 RoBERTa
V-A7 XLNet
V-A8 ProphetNet
V-A9 Falcon
V-A10 Reformer
V-A11 PaLM
V-A12 PaLM2
V-A13 LLaMA
V-A14 LLaMA2
V-A15 GShard
V-A16 ELECTRA
V-A17 MPT-30B
V-A18 Yi-34B
V-A19 Falcon2-11B
V-B LLMs performance in the hardware cybersecurity
Table IV compares 19 publicly available LLMs’ performance in Hardware design-related bug detection and security issues identification using samples from various sources. A portion of the Chrystalis dataset [87] has been used to evaluate the performance of the LLM models in bug detection tasks. A set of faults has been injected intentionally into a functional code and labeled as faulty. The sample size that has been processed comprises 10K of hardware design-related code samples. The prompt that has been used instructs the model to check the concerned code for any issue or bug and respond only with yes or no. The result is presented as the ratio of the responses where the model successfully identified a buggy code from the total samples used. The hardware CWE column evaluates the capability of the models to link a buggy code to its CWE number. The prompt has been designed to ask for a well-defined CWE number on the buggy design. This evaluation process asses the capability of an LLM model in bug detection and classification into the correct CWE class number.
@@ V-C LLMs performance in the cybersecurity knowledge
Table V compares various 42 LLMs performance in the cybersecurity domain using CyberMetric dataset [89]. Figure 5 presents the LLMs performance steps. The models are evaluated based on their accuracy across four question sets: 80 questions, 500 questions, 2000 questions, and 10,000 questions. The performance is represented in percentage accuracy, offering a comprehensive view of each model’s proficiency in handling cybersecurity-related queries.
@@ VI Code-specific LLMs
The rapid evolution of technology and software development has increased the demand for specialized tools that aid in coding, debugging, and enhancing software security [115, 116]. Recognizing this need, various organizations have developed Code-specific LLMs, each offering unique features and capabilities. These models leverage advanced machine learning techniques to understand, generate, and manipulate code, thereby revolutionizing the field of software development [117, 118]. This section delves into several notable Code-specific LLMs, exploring their architectures, training methods, and potential applications in cybersecurity and beyond [119, 120, 121, 122]. Table IX and Table X compare Code-specific Large Language Models.
@@ VI-A9 DeciCoder-1B
TABLE VI: Comparison of Large Language Models
TABLE VII: Continued
TABLE VIII: Continued
TABLE IX: Comparison of Code-specific Large Language Models
TABLE X: Continued
TABLE XI: Datasets Used for Pre-training Foundation Models in Coding
@@ VI-B Datasets Development for Code-centric LLM Models
The development of large-scale datasets has played a crucial role in advancing LLM models, especially those focused on understanding and generating code. Table XI presents the datasets used for pre-training foundation models in Coding. Datasets like CodeSearchNet [145] and The Pile [146] have been instrumental in bridging the gap between natural language and code, improving semantic search capabilities, and enhancing language model training across diverse domains. These datasets provide a rich source of real-world code in multiple programming languages and include expert annotations and natural language queries that challenge and push the boundaries of LLM performance in code-related tasks.
@@
TABLE XII: Comparative Analysis of Vulnerabilities in LLM-Generated Code
@@ VI-C Vulnerabilities Analysis of LLM-Generated Code
The evolution of LLMs in software development has brought significant advancements and new security challenges [158]. Table XII presents a comparative analysis of vulnerabilities in LLM-generated code.
@@ VII-B Software Cyber Security datasets
In software cyber security, datasets play a crucial role in understanding, detecting, and mitigating vulnerabilities in software systems. This sub-section explores several significant software cybersecurity datasets, each offering unique insights and methodologies for vulnerability analysis in cybersecurity. From the extensive BigVul dataset, which links vulnerabilities in the CVE database to specific code commits, to the innovative FormAI dataset, leveraging AI-generated C programs and advanced verification methods for precise vulnerability classification, each dataset contributes uniquely to the field. These datasets range from manually labeled by security experts to those generated using state-of-the-art automated tools, providing diverse resources for researchers and practitioners. Table XIII provides an overview of software vulnerability datasets that can be used for fine-tuning LLMs for software security.
@@ VII-B9 CrossVul
TABLE XIII: Overview of Software Vulnerability Datasets that can be used for fine-tuning LLMs for software security.
@@ VIII LLM vulnerabilities and Mitigation
This section reviews the OWASP Top 10 for LLM Applications project [175], a comprehensive initiative designed to increase awareness about LLM security vulnerabilities. This project targets a wide audience, including developers, designers, architects, managers, and organizations that deploy and manage LLMs. Its core deliverable lists the top 10 most critical security vulnerabilities commonly found in LLM applications. In addition, we include other LLM vulnerabilities not included in the OWASP project, as presented in Table XIV and 7.
@@ VIII-B3 Potential Attack Scenarios
TABLE XIV: Overview of LLM vulnerabilities and Mitigation
@@ VIII-C Adversarial Natural Language Instructions
Wu et al. [184] proposed presented DeceptPrompt, highlighting a critical vulnerability in Code LLMs: their susceptibility to adversarial natural language instructions. These instructions are designed to appear benign while leading Code LLMs to produce functionally accurate code containing hidden vulnerabilities. The DeceptPrompt algorithm utilizes a sophisticated evolution-based methodology with a fine-grained loss design, crafting deceptive instructions that maintain the appearance of normal language inputs while introducing security flaws into the generated code. This vulnerability is exacerbated by the challenges in preserving the code’s functionality, targeting specific vulnerabilities, and maintaining the semantics of the natural language prompts.
@@ VIII-D Automatic adversarial prompt generation
In the highlighted study, Zou et al. [185] address the challenge of automatic adversarial prompt generation in aligned language models. Their method focuses on crafting a specific suffix that maximizes the likelihood of generating objectionable content when attached to various queries directed at an LLM. This approach differs significantly from previous methods by employing automated techniques, including greedy and gradient-based search strategies. The novelty lies in the method’s ability to bypass the alignment measures designed to prevent undesirable outputs from LLMs, effectively circumventing the safeguards put in place. Therefore, the study’s findings underscore the need for robust prevention strategies against these sophisticated adversarial attacks. Enhanced security measures are required to safeguard aligned LLMs. These could include developing more advanced alignment algorithms resistant to adversarial manipulations and implementing real-time monitoring systems to detect and neutralize such attacks. In addition, continuously updating and training models with new adversarial examples can help build resilience against these evolving threats. The focus should be on creating systems that can quickly adapt and respond to the changing tactics of adversarial prompt generation.
@@ VIII-H3 Prevention Strategies
TABLE XV: LLM Cybersecurity Insights.
@@ IX-B LLM Cybersecurity Insights
Table XV presents various facets of LLM integration into cybersecurity, providing insights into architectural nuances, dataset creation, pre-training, fine-tuning methodologies, evaluation metrics, advanced techniques, deployment strategies, security measures, and optimization approaches.
@@ IX-B4 Supervised Fine-Tuning
TABLE XVI: Comparison of Benchmarks for Evaluating LLMs in Cybersecurity Knowledge
@@ IX-B5 Cyber Security Evaluation
Table XVI compares benchmarks for evaluating LLMs in cybersecurity knowledge. CyberMetric [89] is a benchmark dataset designed explicitly for evaluating large language models in knowledge cybersecurity. It consists of 10,000 questions derived from various authoritative sources within the field. The dataset is used to measure the knowledge of LLMs across a spectrum of cybersecurity topics, facilitating direct comparisons between human expertise and machine capabilities. This unique dataset aids in understanding the strengths and limitations of LLMs in cybersecurity, providing a foundation for further development and specialized training of these models in this critical area.
@@
TABLE XVII: Optimization Strategies for Large Language Models in Cybersecurity
@@ IX-B9 Optimizing LLMs
Optimizing LLMs for production encompasses several crucial techniques to enhance speed, reduce memory requirements, and maintain output quality. One pivotal strategy is model quantization, which significantly reduces the precision of model weights—often to 4-bit or 8-bit—thereby decreasing the GPU memory requirements. Table XVII presents the optimization strategies for LLMs that can be adopted for Cybersecurity use cases. For instance, quantizing a model to 4-bit can bring down the VRAM requirement from 32 GB to just over 9 GB, allowing these models to run efficiently on consumer-level hardware like the RTX 3090 GPU. Therefore, advanced attention mechanisms such as Flash Attention reduce computation times by optimizing self-attention layers, which are integral to transformers [219]. This optimization is especially beneficial for handling long input sequences, where traditional self-attention mechanisms could become prohibitively expensive regarding memory and processing power [235, 236].

### v2

@@ V-C LLMs performance in the cybersecurity knowledge
Table IX compares various 42 LLMs performance in the cybersecurity domain using CyberMetric dataset [104] . Figure 5 presents the LLMs performance steps. The models are evaluated based on their accuracy across four question sets: 80 questions, 500 questions, 2000 questions, and 10,000 questions. The performance is represented in percentage accuracy, offering a comprehensive view of each model’s proficiency in handling cybersecurity-related queries.
@@
TABLE X: Comparison of Code-specific Large Language Models
TABLE XI: Continued
@@ VI Code-specific LLMs
The rapid evolution of technology and software development has increased the demand for specialized tools that aid in coding, debugging, and enhancing software security [152, 153]. Recognizing this need, various organizations have developed Code-specific LLMs, each offering unique features and capabilities. These models leverage advanced machine learning techniques to understand, generate, and manipulate code, thereby revolutionizing the field of software development [154, 155]. This section delves into several notable Code-specific LLMs, exploring their architectures, training methods, and potential applications in cybersecurity and beyond [156, 157, 158, 159]. Table X and Table XI compare Code-specific Large Language Models.
@@ VI-A9 DeciCoder-1B
TABLE XII: Datasets Used for Pre-training Foundation Models in Coding
@@ VI-B Datasets Development for Code-centric LLM Models
The development of large-scale datasets has played a crucial role in advancing LLM models, especially those focused on understanding and generating code. Table XII presents the datasets used for pre-training foundation models in Coding. Datasets like CodeSearchNet [160] and The Pile [161] have been instrumental in bridging the gap between natural language and code, improving semantic search capabilities, and enhancing language model training across diverse domains. These datasets provide a rich source of real-world code in multiple programming languages and include expert annotations and natural language queries that challenge and push the boundaries of LLM performance in code-related tasks.
@@
TABLE XIII: Comparative Analysis of Vulnerabilities in LLM-Generated Code
@@ VI-C Vulnerabilities Analysis of LLM-Generated Code
The evolution of LLMs in software development has brought significant advancements and new security challenges [173]. Table XIII presents a comparative analysis of vulnerabilities in LLM-generated code.
@@ VII-B Software Cyber Security datasets
In software cyber security, datasets play a crucial role in understanding, detecting, and mitigating vulnerabilities in software systems. This sub-section explores several significant software cybersecurity datasets, each offering unique insights and methodologies for vulnerability analysis in cybersecurity. From the extensive BigVul dataset, which links vulnerabilities in the CVE database to specific code commits, to the innovative FormAI dataset, leveraging AI-generated C programs and advanced verification methods for precise vulnerability classification, each dataset contributes uniquely to the field. These datasets range from manually labeled by security experts to those generated using state-of-the-art automated tools, providing diverse resources for researchers and practitioners. Table XIV provides an overview of software vulnerability datasets that can be used for fine-tuning LLMs for software security.
@@ VII-B9 CrossVul
TABLE XIV: Overview of Software Vulnerability Datasets that can be used for fine-tuning LLMs for software security.
@@ VIII LLM vulnerabilities and Mitigation
This section reviews the OWASP Top 10 for LLM Applications project [190], a comprehensive initiative designed to increase awareness about LLM security vulnerabilities. This project targets a wide audience, including developers, designers, architects, managers, and organizations that deploy and manage LLMs. Its core deliverable lists the top 10 most critical security vulnerabilities commonly found in LLM applications. In addition, we include other LLM vulnerabilities not included in the OWASP project, as presented in Table XV. Figure 7 presents LLM vulnerabilities included in the OWASP project.
@@ VIII-B3 Potential Attack Scenarios
TABLE XV: Overview of LLM vulnerabilities and Mitigation
@@ VIII-C Adversarial Natural Language Instructions
Wu et al. [199] proposed presented DeceptPrompt, highlighting a critical vulnerability in Code LLMs: their susceptibility to adversarial natural language instructions. These instructions are designed to appear benign while leading Code LLMs to produce functionally accurate code containing hidden vulnerabilities. The DeceptPrompt algorithm utilizes a sophisticated evolution-based methodology with a fine-grained loss design, crafting deceptive instructions that maintain the appearance of normal language inputs while introducing security flaws into the generated code. This vulnerability is exacerbated by the challenges in preserving the code’s functionality, targeting specific vulnerabilities, and maintaining the semantics of the natural language prompts [200].
@@ VIII-D Automatic Adversarial Prompt Generation
VIII-D1 Nature of the Attack
Zou et al. [201] propose a method for automatically generating adversarial prompts in aligned language models. Specifically, they craft a targeted suffix that, when appended to diverse LLM queries, maximizes the likelihood of producing objectionable or undesirable content. Unlike earlier approaches, this method employs automated techniques such as greedy and gradient-based search to circumvent existing alignment measures systematically. By exploiting gaps in the alignment framework, attackers can bypass safeguards to prevent harmful or prohibited outputs.
VIII-D2 Prevention Strategies
The findings from Zou et al. [201] highlight the need for comprehensive defenses against automated adversarial prompt generation. Potential countermeasures include:
  • Advanced Alignment Algorithms: Develop new alignment strategies that are more resistant to adversarial manipulations, ensuring that maliciously crafted suffixes cannot easily override guardrails.
  • Real-Time Monitoring: Implement systems capable of detecting suspicious prompt patterns or gradients in real time, enabling swift neutralization of emerging attacks.
  • Ongoing Model Retraining: Continuously update models with fresh adversarial examples to bolster resilience against newly discovered attack vectors.
  • Adaptive Response Mechanisms: Design LLMs and supporting infrastructure to adapt to changing tactics, reducing the window of opportunity for adversaries.
VIII-D3 Potential Attack Scenarios
Automated adversarial prompt generation can pose significant risks in various contexts:
  • Widespread Content Manipulation: Attackers may propagate malicious suffixes across large-scale user interactions (such as web forums or social media), causing misaligned outputs on a broad scale.
  • Targeted Model Evasion: In specialized applications like content filtering or customer support bots, adversaries might exploit gradient-based techniques to bypass specific policy checks repeatedly.
  • Dynamic Model Attacks: As new alignment protocols are introduced, attackers can use automated search methods to uncover fresh vulnerabilities, fueling an ongoing arms race between defenders and adversaries.
@@ VIII-H3 Prevention Strategies
TABLE XVI: LLM Cybersecurity Insights.
@@ IX-B LLM Cybersecurity Insights
Table XVI presents various facets of LLM integration into cybersecurity, providing insights into architectural nuances, dataset creation, pre-training, fine-tuning methodologies, evaluation metrics, advanced techniques, deployment strategies, security measures, and optimization approaches.
@@ IX-B4 Supervised Fine-Tuning
TABLE XVII: Comparison of Benchmarks for Evaluating LLMs in Cybersecurity Knowledge
@@ IX-B5 Cyber Security Evaluation
Table XVII compares benchmarks for evaluating LLMs in cybersecurity knowledge. CyberMetric [104] is a benchmark dataset designed explicitly for evaluating large language models in knowledge cybersecurity. It consists of 10,000 questions derived from various authoritative sources within the field. The dataset is used to measure the knowledge of LLMs across a spectrum of cybersecurity topics, facilitating direct comparisons between human expertise and machine capabilities. This unique dataset aids in understanding the strengths and limitations of LLMs in cybersecurity, providing a foundation for further development and specialized training of these models in this critical area.
@@
TABLE XVIII: Optimization Strategies for Large Language Models in Cybersecurity
@@ IX-B9 Optimizing LLMs
Optimizing LLMs for production encompasses several crucial techniques to enhance speed, reduce memory requirements, and maintain output quality. One pivotal strategy is model quantization, which significantly reduces the precision of model weights—often to 4-bit or 8-bit—thereby decreasing the GPU memory requirements. Table XVIII presents the optimization strategies for LLMs that can be adopted for Cybersecurity use cases. For instance, quantizing a model to 4-bit can bring down the VRAM requirement from 32 GB to just over 9 GB, allowing these models to run efficiently on consumer-level hardware like the RTX 3090 GPU. Therefore, advanced attention mechanisms such as Flash Attention reduce computation times by optimizing self-attention layers, which are integral to transformers [235]. This optimization is especially beneficial for handling long input sequences, where traditional self-attention mechanisms could become prohibitively expensive regarding memory and processing power [251, 252].